# RLSA-PFL: Robust Lightweight Secure Aggregation with Model Inconsistency Detection in Privacy-Preserving Federated Learning

### v2

@@ 7.2 Experimental Result
We developed a prototype of our scheme and tested it on a VM with 2 vCPUs (Intel Xeon), 16GB of RAM, and Ubuntu Server 22.04. The prototype is implemented in Python (Python 3.7.11) using PyTorch 1.13.1444https://pytorch.org/, NumPy 1.21.6555https://pypi.org/project/numpy/1.21.6/, and PyCryptodome 3.20666https://pypi.org/project/pycryptodome/.
7.2.1 Running Time Analysis Without User Dropout
Figure 4 and Figure 6 present the running time of an intermediate server and the aggregator during a round without any user dropouts in the Semi-Honest Model, with varying numbers of users and input vector sizes, respectively. Similarly, Figure 5 and Figure 7 show the running time for the same scenario in the Malicious Model. We considered randomly generated input vectors with $64$-bit entries as the users’ local gradients. In this analysis, the average running time of one intermediate server was measured, as all intermediate servers operate in parallel. The results indicate that as the number of users and input vector sizes increase, the average running time per round grows linearly, causing the total running time for both the server and intermediate servers to scale linearly with the number of users and input vector size.
7.2.2 Running Time Analysis With User Dropout
In the next sets of experiments, we consider user dropouts to measure the total running time of the user, intermediate server, and aggregator, along with capturing the running time of each main phase of our scheme. Table II and Table III present the running times (in microseconds) per round for users, intermediate servers, and the aggregator during various phases under the semi-honest and malicious models, respectively. It compares scenarios with different percentages of user dropouts ($0\%,10\%,20\%,30\%$) and varying user counts ($500$ and $1000$). For both models, the experiments used $50k$ random input vectors with $64$-bit entries and $10$ intermediate nodes.
For each scenario, the masking time at the user remains relatively stable, showing minimal variation as the percentage of user dropouts and the number of users increase, as it is independent of both. However, the partial aggregation time at the intermediate server and final aggregation time at the aggregator increase with the number of users. The running time decreases with more user dropouts due to a reduction in the number of active participants. The global model verification time at both the aggregator and user remains consistent across different scenarios, independent of user participation and dropouts. The total running times for the aggregator and intermediate server also increase with the number of users but decrease as user dropouts increase. The total running time for the user remains stable as it is unaffected by the number of participants or dropouts.
8 Conclusion
In this paper, we introduced a lightweight secure aggregation scheme for FL that uses a key negation-based masking technique. Our approach adds an intermediate server layer, which eliminates the need for users to communicate with each other and only requires one setup phase. This significantly reduces the burden on users. Thanks to this intermediate layer, our scheme can handle user dropouts effectively, allowing users to join the training phase at any time with just one additional round of communication between the intermediate servers and the aggregator.
Additionally, we integrated a lightweight verification mechanism for model parameters to detect inconsistencies, enhancing security against malicious aggregators. We conducted a thorough formal security analysis showing that our scheme is robust against both semi-honest and malicious environments. Our comparisons with closely related works demonstrate that our scheme is better in terms of security, functionality, and both communication and computation costs. We also implemented our scheme and presented comprehensive experimental results, proving its practical usability in real-world applications.
@@ 8 Conclusion
Appendix A Security Proof Against Semi-Honest Aggregators and Intermediate Servers
Proof 13
@@ Appendix A Security Proof Against Semi-Honest Aggregators and Intermediate Servers
This hybrid is similar to the previous one. The only difference is that the simulator $\mathsf{Sim}$ replaces the random masked $r_{i}$ for each users in $(\mathbb{U}_{1}\setminus C_{c})$ with $x_{i}$ subject to
$\sum_{i\in\mathbb{U}_{1}\setminus C_{c}}r_{i}=$ $\sum_{\mathbb{U}_{1}\setminus C_{c}}x_{i}$ (9)
As we can observe the distribution of the masking values generated based on the symmetric homomorphic encryption scheme for masking $r_{i}$ and $x_{i}$ is identical subject to the Eq. 9.
Therefore, the last hybrid shows that we can define a simulator $\mathsf{Sim}$ which is computationally indistinguishable from that of the real execution of $\mathsf{Real}$ from the combined views of the honest-but-curious entities. Hence, it completes the proof.
Appendix B Security Proof Against Dishonest Users, Intermediate Servers and Aggregator
Proof 14
We use the standard hybrid argument to carry out the security proof of this theorem. Similar to the Theorem 5, we construct a simulator $\mathsf{Sim}$ by the subsequent modifications to the random variable $\mathsf{Real}$. Our goal is to prove that two subsequent hybrids are computationally indistinguishable to ensure that the distribution of simulator $\mathsf{Sim}$ as a whole is also identical to the real execution $\mathsf{Real}$.
@@ Appendix B Security Proof Against Dishonest Users, Intermediate Servers and Aggregator
In this hybrid scenario, the simulator $\mathsf{Sim}$ replaces the masked locally trained model $x_{i}$ of each honest user $u_{i}$ with a randomly generated number, while it replaces it with 00 for the adversaries. Based on the security of the pseudo-random function, the joint views of the users in $M_{c}$ are identical to those of $\mathsf{Real}$. Hence, this hybrid is indistinguishable from the previous one.
@@
In this hybrid, the simulator $\mathsf{SIM}$ queries to the Random Oracle $\mathcal{O}$ for the values $w_{i}$ for the set $(\mathrm{I}\setminus C_{c})$ with respect to $\sum_{u_{i}\in\mathcal{Q}\setminus C_{c}}w_{i}=\sum_{u_{i}\in\mathcal{Q}\setminus C_{c}}x_{i}$ instead of receiving the inputs of the honest users and masks it with a random number. For the adversaries, the $\mathsf{Sim}$ chooses random numbers (masks are set to be 00 here). The Random Oracle $\mathcal{O}$ will not abort and does not modify the joint view of $M_{c}$ if at least more than $2$ participants are honest due to the properties of the random oracle and pseudo-random function. Thus, this hybrid is indistinguishable from the previous one.
@@
It can be observed from the last hybrid that the adversary’s view remains unaffected which proves the indistinguishable property between the different hybrids. Furthermore, this hybrid does not take input from honest parties. This completes the proof.

### v3

@@ 8. Experimental Evaluation
We developed a prototype of our protocol and tested it on a VM with 2 vCPUs (Intel Xeon), 16GB of RAM, and Ubuntu Server 22.04. The prototype is implemented in Python (Python 3.7.11) using PyTorch 1.13.1, NumPy 1.21.6, and PyCryptodome 3.20. Our prototype employs the LeNet-5 architecture as an example to demonstrate both the learning performance and practical applicability of our protocol.
For performance testing, we omitted the actual model training process and instead generated a random data vector of the required size. This allowed us to obtain performance results comparable to those of e-SeaFL and to observe the scalability trend as the weight size increased. The implementation used for performance testing of our protocol is available at https://anonymous.4open.science/r/scatesfl-815A/. The real-world federated model training (using MNIST or CIFAR-10) based on our protocol can be found at https://anonymous.4open.science/r/scatesfl-44E1/. The corresponding e-SeaFL implementation is accessible at https://anonymous.4open.science/r/scatesfl-6F59/.
(a) MNIST
(b) CIFAR-10
Figure 8. Accuracy Comparison Between Our Protocol and Baseline FedAvg (McMahan et al., 2017)
8.1. User Running Time Computation
Table 3 presents the average per-user computation time across different phases- masking, model verification, and total running time- using a vector size of $48k$ with 64-bit elements and 5 intermediate servers per round. These values are averaged over five rounds and remain nearly constant across varying numbers of users, as the computation cost is primarily influenced by the input vector size.
From Table 3, it is evident that our protocol reduces masking time by about $98\%$ in the semi-honest model and $97\%$ in the malicious model, and reduces model verification time by about $99\%$ in both models, compared to e-SeaFL. As discussed earlier, e-SeaFL incurs higher computation due to additional PRF operations in the masking phase and elliptic curve cryptographic operations during model verification. Moreover, the reported masking time for e-SeaFL in our comparison excludes the computations related to the commitment phase (please refer e-SeaFL (Behnia et al., 2024) for more details on the commitment phase); when these are included, the masking time increases to approximately $1.6$ minutes. Even without considering the commitment phase, our protocol demonstrates clear computational advantages, and when it is included, the performance gap becomes even more pronounced— underscoring our protocol’s efficiency in both phases.
[…]
8.2. Intermediate Server and Aggregator Running Time Computation
Figures 4(a) and 4(b) illustrate the computation times of the intermediate server and aggregator, respectively, under the semi-honest model, comparing our protocol with e-SeaFL across vector sizes of 16K, 32K, and 48K, and user counts up to 500. Experimental results show that our protocol reduces intermediate server computation time by approximately 600 to 2100 times, and aggregator computation time by approximately 300 to 580 times, compared to e-SeaFL. Figures 5(a) and 5(b) present results under the malicious model, where additional signature operations are required. Our experiments show that, compared to e-SeaFL, our protocol reduces intermediate server computation time by approximately $60$ to $165$ times, and aggregator computation time by approximately $220$ to $510$ times.
We further evaluated performance by varying the vector size (16K to 64K) for 100, 200, and 500 users. Figures 6(b) and 7(b) show the average computation time of the aggregator in the semi-honest and malicious models, respectively. Similarly, Figures 6(a) and 7(a) show the average computation time of the intermediate server in the semi-honest and malicious models, respectively. From these figures, it is evident that our protocol reduces aggregator computation time by approximately $250$ to $560$ times in the semi-honest model and $220$ to $510$ times in the malicious model compared to e-SeaFL. Similarly, it reduces intermediate server computation time by approximately $350$ to $1900$ times in the semi-honest model and $47$ to $154$ times in the malicious model.
8.2.1. Accuracy Comparison
We conducted experiments on two benchmark datasets, CIFAR-10 (Krizhevsky et al., 2009) and MNIST (Lecun et al., 1998), to assess the learning performance. CIFAR-10 has 60,000 color images of size $32\times 32$ pixels, divided into 10 different classes. It has 50,000 training images and 10,000 test images. This dataset is used to test models on more challenging image tasks. MNIST has 70,000 grayscale images of handwritten digits, each $28\times 28$ pixels. It includes 60,000 training images and 10,000 test images. MNIST is commonly used to check how well models learn simple image recognition tasks quickly.
Figures 8(a) and 8(b) present the accuracy of our protocol during training compared to the baseline FedAvg algorithm (McMahan et al., 2017), over 500 rounds with 100 users. While FedAvg achieves slightly higher final accuracy, our protocol demonstrates steady and reliable learning throughout the process. On CIFAR-10, our protocol reaches 38% accuracy versus FedAvg’s 45%, and on MNIST it attains 89.9%, which is close to FedAvg’s 96%. The small drop in accuracy comes from encoding floating-point values as integers, where a minor loss of precision accumulates across rounds. Despite this, our protocol maintains strong accuracy while providing additional benefits in privacy, communication efficiency, and reliability, making it a practical and effective choice in settings with system constraints or strict security requirements.
9. Conclusion
In this paper, we presented a novel secure aggregation protocol tailored for dynamic FL environments. Our design addresses key limitations of existing protocols by introducing a masking protocol based on additive symmetric homomorphic encryption with key negation. This approach eliminates the need for repeated cryptographic setup, significantly reducing communication and computation overhead. Our protocol is the first single-setup protocol to simultaneously achieve forward and backward secrecy, model integrity verification, and dynamic user participation— all while maintaining strong privacy guarantees and resilience to user dropouts. By incorporating intermediate servers, our protocol avoids user-to-user communication, reduces trust assumptions, and enhances scalability for real-world deployments. We also provided formal security proofs under both semi-honest and malicious adversarial models, demonstrating robustness. Our experimental evaluation shows that our protocol achieves up to $99\%$ reduction in user-side computation time compared to state-of-the-art protocols like e-SeaFL, while maintaining competitive model accuracy on benchmark datasets. Overall, our protocol offers a practical, efficient, and secure solution for FL in dynamic and resource-constrained environments, paving the way for broader adoption of privacy-preserving ML.
In our protocol, we assume that the intermediate servers remain online throughout the training process. In certain deployment scenarios, such as ad-hoc or intermittently connected edge networks, this assumption might not hold. Designing a protocol that supports dynamic intermediate servers while maintaining the same functionality, security, and efficiency is an important direction for future work. Additionally, further evaluation on real-world IoT testbeds is planned to validate the protocol’s performance and scalability under practical conditions, with a particular focus on communication overhead.
Acknowledgement
The work has been supported by the Cyber Security Research Centre Limited whose activities are partially funded by the Australian Government’s Cooperative Research Centres Program.
Appendix A Detailed Analysis of Forward and Backward Secrecy in e-SeaFL and Flamingo
Both forward secrecy (FS) and backward secrecy (BS) are critical privacy properties in secure aggregation protocols, ensuring that the compromise of long-term cryptographic material does not expose users’ model updates from prior training rounds and any future training rounds, respectively. This protection is particularly important in FL, where training may span multiple rounds over days, and client devices are often deployed in untrusted or adversarial environments. Without FS and BS, a one-time key compromise can retroactively invalidate the privacy of all prior and future contributions, respectively.
As discussed earlier, single-setup secure aggregation protocols offer a promising direction for scalable, efficient, and secure FL. Compared to multi-setup protocols, they significantly reduce communication and computation overhead by avoiding repeated cryptographic setup in each training round. Two notable protocols in this category are e-SeaFL (Behnia et al., 2024) and Flamingo (Ma et al., 2023), both of which adopt similar masking techniques based on long-term shared secrets and PRFs.
Among the two, e-SeaFL demonstrates improved performance and practicality due to its architectural design, which eliminates direct user-to-user interactions and relies instead on a few assisting nodes for mask coordination and recovery. This makes it more suitable for large-scale FL deployments.
Despite their efficiency, both e-SeaFL and Flamingo share a critical limitation: they do not provide FS and BS. In the following section, we analyze e-SeaFL in detail to demonstrate how its use of long-term cryptographic material exposes it to FS and BS vulnerabilities. Since Flamingo employs a similar masking strategy, the same analysis applies, indicating that Flamingo also lacks both FS and BS.
A.1. Mask Derivation in e-SeaFL
e-SeaFL is a single-setup secure aggregation protocol proposed by Behnia et al. (ACSAC 2024) to improve communication efficiency by avoiding repeated cryptographic setup across rounds. In the setup phase, each user $i$ and assisting node $a\in\mathcal{A}_{i}$ perform a key agreement protocol ($\mathsf{KA}$) to derive a long-term shared secret or seed:
$s_{i,a}=\mathsf{KA}(SK_{i},PK_{a}),$
where $SK_{i}$ and $PK_{a}$ are the private key of user $i$ and the public key of assisting node $a$, respectively. This shared secret $s_{i,a}$ is then used in each training round $t$ as input to a $\mathtt{PRF}$ to generate a round-specific mask value:
(9) $m_{i,a}^{(t)}$ $=\mathtt{PRF}(s_{i,a},t)$
The final masked update sent by user $i$ in round $t$ is constructed as:
$\tilde{x}_{i}^{(t)}=x_{i}^{(t)}+\sum_{a\in\mathcal{A}_{i}}m_{i,a}^{(t)}.$
where $x_{i}^{(t)}$ is the local model update of the user $i$ for the round $t$.
To recover the true global model update $\theta_{t}$, the server collaborates with assisting nodes to remove the aggregated masks. Each assisting node $a$ computes the sum of its masking contributions across all users, say $\mathcal{U}$, it assists:
$\sum_{i\in\mathcal{U}}m_{i,a}^{(t)}=\sum_{i\in\mathcal{U}}\mathtt{PRF}(s_{i,a},t)$
After receiving all the aggregated masking contributions from the assisting nodes, the central server removes these masks from the aggregated global model parameters to recover the plaintext model update.
A.2. Forward and Backward Secrecy Limitations
In e-SeaFL, round-specific masking values $m_{i,a}^{(t)}$ are deterministically derived from a long-term shared secret $s_{i,a}$ established between user $i$ and assisting node $a$. The masking function is defined in the Eq. 9. This design eliminates the need for repeated setup phases, but it introduces a critical vulnerability: if the shared secret $s_{i,a}$ is compromised, an adversary can compute all past (and future as well) masking values for every round. Consequently, the attacker can reconstruct the original model update $x_{i}^{(t)}$ for any round $t$ by subtracting the derived masks from the masked update $\tilde{x}_{i}^{(t)}$:
$x_{i}^{(t)}=\tilde{x}_{i}^{(t)}-\sum_{a\in\mathcal{A}_{i}}m_{i,a}^{(t)},$
where $\mathcal{A}_{i}$ is the set of assisting nodes for user $i$.
A.3. Practical Implications
The absence of forward and backward secrecy in e-SeaFL and Flamingo present a tangible risk in real-world FL deployments. Devices such as smartphones, edge nodes, and IoT sensors often operate in environments with limited physical and software security, making them vulnerable to malware, unauthorized access, and insider threats- especially over prolonged training sessions.
Once a long-term secret $s_{i,a}$ is established between a user and an assisting node, it remains unchanged throughout the training process. If compromised at any point, an adversary can retroactively and proactively derive all masking values, enabling reconstruction of the user’s full update history. This undermines both privacy and the integrity of the aggregation process. While e-SeaFL and Flamingo offer efficiency through its single-setup design, its reliance on static secrets could make it unsuitable for deployments requiring strong long-term privacy guarantees.
Appendix B $\alpha$-summation ideal functionality
As discussed in (Bell et al., 2020) and (Behnia et al., 2024), a summation protocol is referred to as $\alpha$-secure if it ensures that each honest participant’s input is aggregated with at least $\alpha\cdot|\mathcal{H}|$ other secret values, where $\mathcal{H}$ denotes the set of honest parties. We adopt this notion of $\alpha$-security in our simulation-based proof framework, consistent with the above works.
Definition B.1 ($\alpha$-summation ideal functionality).
Let $p,n,d\in\mathbb{Z}$, and let $\alpha\in[0,1]$. Suppose we have a subset $S\subseteq[|\mathcal{U}|]$, and define the collection $\mathbf{X}_{S}=\{\mathbf{x}_{i}\}_{i\in S}$, where each vector $\mathbf{x}_{i}\in\mathbb{Z}_{p}$. Let $\mathcal{P}_{S}$ represent the set of all partitions of $S$, and consider a partition $\{S_{1},\ldots,S_{k}\}\in\mathcal{P}_{S}$ composed of pairwise disjoint subsets.
The $\alpha$-summation ideal functionality is modeled by a function $\mathcal{F}_{\mathbf{X},\alpha}(\cdot)$, which takes as input the partition $\{S_{j}\}_{j=1}^{k}$ and outputs a collection $\{\mathbf{s}_{j}\}_{j=1}^{k}$, where each $\mathbf{s}_{j}$ is computed as follows:
$\mathbf{s}_{j}=\begin{cases}\sum\limits_{i\in S_{j}}\mathbf{x}_{i}&\text{if }|\mathcal{P}_{S}|\geq\alpha\cdot|S|\\ \bot&\text{otherwise}\end{cases}\quad\text{for all }j\in[1,\ldots,k].$
Appendix C Security Against Semi-Honest Users Only
Proof.
In our protocol, users randomly choose their input values (i.e., random secret keys) to mask the local model parameters, similar to a one-time pad. Each user’s input values are independent of those of other users. Since the aggregator and intermediate servers are considered honest entities, the combined view of the users in $C_{c}$ will not reveal any useful information about the input values of users not in $C_{c}$. Additionally, the honest-but-curious users in $C_{c}$ only receive the identities of the honest users not in $C_{c}$. This enables the simulator $\mathsf{Sim_{C}^{\mathbb{U},t_{c},\lambda}}$ to use dummy input values for the honest users not in $C_{c}$ while keeping the combined views of the users in $C_{c}$ identical to that of $\mathsf{Real_{C}^{\mathbb{U},t_{c},\lambda}}$. ∎
@@ Appendix C Security Against Semi-Honest Users Only
Appendix D Security Against Semi-Honest Intermediate Servers Only
Proof.
Appendix E Security Proof Against Semi-Honest Aggregators and Intermediate Servers
Proof.
@@ Appendix E Security Proof Against Semi-Honest Aggregators and Intermediate Servers
This hybrid is similar to the previous one, with the only difference being that the simulator $\mathsf{Sim}$ replaces the random masked values $r_{i}$ for each user in $(\mathbb{U}_{1}\setminus C_{c})$ with values $x_{i}$, subject to the constraint:
(10) $\sum_{i\in\mathbb{U}_{1}\setminus C_{c}}r_{i}=\sum_{i\in\mathbb{U}_{1}\setminus C_{c}}x_{i}$
The distribution of the masking values $r_{i}$ and $x_{i}$, when generated using a symmetric homomorphic encryption scheme, is identical subject to the Eq. 10.
The last hybrid shows that we can define a simulator $\mathsf{Sim}$ which is computationally indistinguishable from that of the real execution of $\mathsf{Real}$ from the combined views of the honest-but-curious entities. Hence, it completes the proof. ∎
Appendix F Security Against Dishonest Intermediate Servers Only
Proof.
This proof is similar to that of Theorem 7.2, as the intermediate servers in $C_{f}$ gain no knowledge about $x_{\mathbb{F}\setminus C_{f}}$ apart from the list of participants. Consequently, the simulator $\mathsf{Sim}$ can assign real inputs to the dishonest intermediate servers while assigning dummy inputs to the remaining intermediate servers, accurately replicating the perspectives of the intermediate servers in $C_{f}$. Thus, the joint views of the intermediate servers in $C_{f}$ in the simulation are indistinguishable from those in $\mathsf{Real}$. ∎
Appendix G Security Proof Against Dishonest Users, Intermediate Servers and Aggregator
Proof.
We use the standard hybrid argument to carry out the security proof of this theorem. Similar to the Theorem 7.3, we construct a simulator $\mathsf{Sim}$ by the subsequent modifications to the random variable $\mathsf{Real}$. Our goal is to prove that two subsequent hybrids are computationally indistinguishable to ensure that the distribution of simulator $\mathsf{Sim}$ as a whole is also identical to the real execution $\mathsf{Real}$.
@@ Appendix G Security Proof Against Dishonest Users, Intermediate Servers and Aggregator
In this hybrid scenario, the simulator $\mathsf{Sim}$ replaces the masked locally trained model $x_{i}$ of each honest user $u_{i}$ with a randomly generated number, while it replaces it with 0 for the adversaries. Based on the security of the pseudo-random function, the joint views of the users in $M_{c}$ are identical to those of $\mathsf{Real}$. Hence, this hybrid is indistinguishable from the previous one.
@@
In this hybrid, the simulator $\mathsf{SIM}$ queries to the Random Oracle $\mathcal{O}$ for the values $w_{i}$ for the set $(\mathrm{I}\setminus C_{c})$ with respect to $\sum_{u_{i}\in\mathcal{Q}\setminus C_{c}}w_{i}=\sum_{u_{i}\in\mathcal{Q}\setminus C_{c}}x_{i}$ instead of receiving the inputs of the honest users and masks it with a random number. For the adversaries, the $\mathsf{Sim}$ chooses random numbers (masks are set to be 0 here). The Random Oracle $\mathcal{O}$ will not abort and does not modify the joint view of $M_{c}$ if at least more than $2$ participants are honest due to the properties of the random oracle and pseudo-random function. Thus, this hybrid is indistinguishable from the previous one.
@@
$\mathsf{Hyb_{10}}$
In this hybrid, the simulator $\mathsf{Sim}$ selects a random secret from $\mathbb{Z}_{q}$ to mask the aggregated model, as in the previous hybrid, and computes a message authentication code (MAC) over the aggregated model using a secret key unknown to the adversary. Since the random masking value is never reused, it effectively functions as a one-time pad. Furthermore, the MAC, assumed to be pseudorandom under the PRF assumption, reveals no information to the adversary. Consequently, this hybrid is computationally indistinguishable from the previous one.
It can be observed from the last hybrid that the adversary’s view remains unaffected which proves the indistinguishable property between the different hybrids. Furthermore, this hybrid does not take input from honest parties. This completes the proof. ∎